OpenStudy (anonymous):

4sinx+3cosx=0 the range is from 0-360 Solve in exact form without using the calculator

OpenStudy (anonymous):

Sorry that's all my teacher specified

OpenStudy (cwrw238):

i think we can do this by dividing through by cos x 4sinx + 3 cosx = 0 ---- ---- cos x cos x 4 tan x + 3 = 0 tan x = -3/4

OpenStudy (anonymous):

alright I got that same answer (: thanks

OpenStudy (cwrw238):

yea but how do you get value of x without a calculator??

OpenStudy (cwrw238):

your welcome

OpenStudy (anonymous):

it was a non calculator question. like that would've been the answer

OpenStudy (anonymous):

take the inverse tangent next, and you'll get x

OpenStudy (cwrw238):

??

OpenStudy (cwrw238):

oh you mean arctan (-3/4)

OpenStudy (anonymous):

it's an ib thing. sorry but that was the answer

OpenStudy (anonymous):

if you wanna know what value of x makes tanx = -3/4, take the inverse

OpenStudy (anonymous):

yeah arctan, i guess that would be more clear to say lol